OpenStudy (anonymous):

if the dimensions of an equilateral triangle are trebled, how is the area affected?

OpenStudy (anonymous):

what is the fastest method?

OpenStudy (anonymous):

\[Area(Equilateral) = \frac{\sqrt{3}}{4} \times a^2\] a has not trebled so a has becomes 3a: \[\implies \frac{\sqrt{3}}{4} \times (3a)^2\] \[\implies 9 \times \frac{\sqrt{3}}{4} \times a^2 \implies 9 \times (Earlier \; Area)\]